kekeman:

What is the sum of the series sum from n equals 1 to infinity of negative 4 times five eighths to the n power question mark negative five halves twenty thirds negative twenty thirds twenty thirteenths

Vocaloid:

\[\sum_{n=1}^{n=\infty}-4(\frac{ 5 }{8 })^{n}\](please double check to make sure I interpreted the series formula correctly) for the sum of an infinite geometric series: \[S_{n}=\frac{ a_{1} }{ (1-r) }\] where a1 is the first term and r is the common ratio. to find a1, plug in n = 1 into the series formula and calculate the first term. for the common ratio r, notice in the series formula, each time n increases by 1, it multiplies the previous term by (5/8) so r = 5/8.

kekeman:

It was negative twenty thirds